Amount of soap used More Happy Less Happy
High 14 86
Moderate 46 54
Low 76 24

1. Calculate the prevalence ratio for the association between moderate amount of soap used
(compared to low amount) and more happiness with life.

2. Based on the table, the conclusion that Soap use causes less happyness with life is: (Select
one)
a. Justified because the risk for being more happy with life is less than 1 for both high and
moderate amount of Soap use compared to low amount of Soap use
b. Justified because the study was conducted among Californians who are representative
of all Soap users
c. Not justified because a cross-sectional
study cannot establish temporality, something that is
necessary in order to establish causality

d. Not justified because the relative risks were weak, thus failing the causal guideline of strength
of association, something that is necessary in order to establish causality
Answer #1

Q. No. 1. Answer :

Prevalance ratio : it is the ratio of the proportion of the persons with disease over the proportion with the exposure.

Formula is here

PR = a/a+ b / c/c+ d

Here a = 46

b = 54

c = 76

d = 24.

So if we put in formula means = 46/46+54 / 76/76+24

= 46/100 / 76/100

= 0.46/0.76

Prevalance ratio is  = 0.6052 .

Q. No. 2. Answer :

Answer is " a "

Justified because the risk for being more happy with life is less than one for both high and moderate amount of soap use compared to low amount of soap use.